Is the mean lifetime of a particular car engine at least 300,000 miles? To test this claim, a sample of 27 engines is measured, yielding an average of 295,450 miles and a standard deviation of 4200 miles. Use a significance level of 0.01 to test this hypothesis.
A) IEETS B) NEETS C) IEETR D) NEETR ​E) None of the Above
Answer #1

Let the population mean lifetime of a particular car engine be \mu

Here we are to test

H_0:\mu\geq 300000\;\;against\;\;H_1:\mu< 300000

The given data is summarized as

Sample size(n)=27

Sample mean(\bar{x})=295450

Sample SD(s)=4200

The test statistic is given by

t=\frac{\bar{x}-300000}{\frac{s}{\sqrt{n}}}

\Rightarrow t=\frac{295450-300000}{\frac{4200}{\sqrt{27}}}=\frac{-4550}{808.2903769}=-5.629165125

The test statistic follows t distribution with df 26

The p-value is obtained to be lesser than 0.00001

As the p-value is less than 0.01, we reject the null hypothesis at 1% level of significance and hence conclude that the population mean lifetime of a particular car engine is less than 300000

The correct option is option C

Hopefully this will help you. In case of any query, do comment. If you are satisfied with the answer, give it a like. Thanks.